# Scalable Offline Metrics for Autonomous Driving

### v1

@@ Abstract
Real-World evaluation of perception-based planning models for robotic systems, such as autonomous vehicles, can be safely and inexpensively conducted offline, ı.e., by computing model prediction error over a pre-collected validation dataset with ground-truth annotations. However, extrapolating from offline model performance to online settings remains a challenge. In these settings, seemingly minor errors can compound and result in test-time infractions or collisions. This relationship is understudied, particularly across diverse closed-loop metrics and complex urban maneuvers. In this work, we revisit this undervalued question in policy evaluation through an extensive set of experiments across diverse conditions and metrics. Based on analysis in simulation, we find an even worse correlation between offline and online settings than reported by prior studies, casting doubts on the validity of current evaluation practices and metrics for driving policies. Next, we bridge the gap between offline and online evaluation. We investigate an offline metric based on epistemic uncertainty, which aims to capture events that are likely to cause errors in closed-loop settings. The resulting metric achieves over 13% improvement in correlation compared to previous offline metrics. We further validate the generalization of our findings beyond the simulation environment in real-world settings, where even greater gains are observed.
@@ I Introduction
Can we estimate the online performance of a driving model without actively driving with it? Evaluating driving policy models that map camera input to a driving plan or controller action in real-world conditions is typically expensive and potentially hazardous. During model deployment, traffic infractions and collisions may occur, such as fatal accidents with pedestrians and vehicles [1, 2], Waymo’s driving in the wrong lane and veering unsafely [3], and Tesla’s Autopilot crashing into a truck on a busy highway [4].
To circumvent such safety-critical evaluations, researchers often use an offline approach. They leverage a computed expected loss over a pre-collected, real-world dataset. This method avoids the potentially dangerous and costly quantification of collisions and traffic infractions in live scenarios (Fig. 1). The efficiency in offline evaluation over a pre-collected driving dataset, where the driving model predictions are compared against safe ground-truth targets provided by a human driver (e.g., leveraging an L2 loss [5, 6, 7, 8, 9, 10, 11, 12, 13]), has recently propelled significant advancements in scalable autonomous driving models [14, 7, 10, 15, 16, 17, 18, 19]. However, it remains unclear whether current offline evaluation practices adequately reflect actual on-road driving model performance, particularly for decisions that require highly robust and split-second reactions with life-threatening implications.
@@
The reasons for the discrepancy between offline and online model performance are three-fold. First, real-world driving datasets are nearly entirely uneventful, with drivers rarely experiencing safety-critical events such as abrupt pedestrian crossings, near-collisions, or even turns and merging. The expected loss over the inherently imbalanced distribution of data can obscure performance under safety-critical events [20, 21]. Second, while automatic data sampling techniques can alleviate data imbalance issues [22, 23, 24, 25, 26], the standard L2 loss-based metric itself provides a crude measure for the ultimate task of safe and enjoyable driving [27, 28]. For instance, the L2 metric can remain identical among several maneuvers, e.g., merging to the left or right lane on a freeway, yet one decision may be perfectly safe while the other results in a collision, i.e., in the case of an adjacent vehicle in a nearby lane. This observation holds for nearly any scenario in complex urban driving settings, as shown in Fig. 1. Additional open-loop metrics have been devised, such as those based on future collision estimates [5, 7, 9, 10, 21]. However, these metrics induce similar evaluation issues. Their reliance on various naive velocity models prevents them from capturing realistic interactive navigation and negotiation in intricate and dense settings. Moreover, collision metrics lack scalability due to the requirement of perception annotations [29]. Third, online model testing is done in a closed-loop manner, where minor decision errors can propagate and result in unpredictable model behavior over time in testing [30, 31, 32, 33].
@@
Despite known limitations in offline open-loop evaluation [19, 21], prior work has rarely studied extrapolation from offline to online settings beyond simplistic settings. For instance, Codevilla et al. [28] demonstrated the limited correlation between the two settings but only in the simplest towns and maneuvers in the CARLA simulation [34], i.e., Town01 and Town02, with one online metric (success rate), and partial action values (only steer, excluding throttle and brake ground truth). Therefore, generalization to intricate maneuvers and complex urban scenarios, diverse online metrics, or real-world settings remains understudied despite its broad implications. In this work, we revisit the question of the correlation between offline and online evaluations.
@@
Contributions: We aim to advance safe and reliable evaluation procedures for autonomous driving policies. Our key insights are threefold. First, we leverage a more expansive evaluation in simulation, covering multiple towns and closed-loops metrics. Our results reveal even worse correlation patterns than have been previously shown. Subsequently, we hypothesize that the poor correlation is due to degraded model performance during instances of uncertainty, i.e., challenging or rare events in the data. We thus analyze an effective metric based on an uncertainty-weighted error, and demonstrate its broad effectiveness across settings, models, and error types. While uncertainty-based techniques are standard in robotics and machine learning [35, 36, 37], to the best of our knowledge, we are the first to quantify their role in estimating online evaluation of driving policies. We observe a significant improvement, e.g., 13% higher correlation in CARLA. Third, given challenges in realistic simulation of real-world physics and appearance, we go beyond simplistic simulations to characterize correlation trends in the real world using a small-scale platform, further validating our proposed metric.
@@ II Related Work
End-to-End Autonomous Driving: With the emergence of deep learning models, autonomous driving research has undergone a significant transformation, shifting from rule-based systems to learning-based approaches. This shift is notably reflected in the proliferation of end-to-end autonomous driving policies utilizing deep learning techniques [38, 39, 40, 41, 8, 42, 14, 43, 6, 44, 17]. This trend is further exemplified by the CARLA Leaderboard [45], where deep learning-based methods have gained prominence, surpassing modular approaches in autonomous driving performance [46, 47, 48, 49, 50, 51]. In addition to camera-based RGB images, additional sensors, such as LiDAR, have also been shown to play a key role in improving the safety of driving policies through multi-modal sensor fusion [41, 52, 53].
@@
Issues with Offline Evaluation: The progress of learning-based models is primarily driven by offline evaluation methods. However, the question of their accuracy in assessing driving performance within online settings still remains. Codevilla et al. [28] is one of the first to attempt to study this problem and proposes several novel offline evaluation metrics, including speed-weighted and quantized versions of standard loss metrics. Among the metrics, the NuScenes Detection Score (NDS) [54] is a widely accepted offline metric that has been shown in a recent work by Schreier et al. [29] to provide improved correlation with online metrics like CARLA Driving Score [34]. As shown by Li et al. in [21], a simple baseline model without vision can achieve competitive status on a leaderboard evaluated by offline metrics like NDS, thus proving the limitations of existing offline metrics in predicting driving performance. These recent discussions demand a revisit on whether offline metrics are sufficient in predicting driving performance.
@@
Safety-Critical Scenario Analysis: Safety critical scenarios refer to specific scenarios, such as collisions, where safety is compromised, and the ego vehicle must take preventative or evasive actions to avoid catastrophic outcomes [4, 55, 25, 56]. Current methodologies address this challenge by curating datasets with generated safety-critical scenarios [26, 57, 23, 22]. This approach allows for testing on various generated scenarios to validate if the policy is safe. While we share a similar goal, this is an orthogonal direction to our approach as we focus on designing a useful metric across datasets. However, while generated scenario realism and coverage remain an open question, our proposed metric can be automatically and easily applied at scale to any policy and dataset (i.e., only based on the model and without requiring privileged annotated information).
@@
Off-Policy Evaluation: Offline Policy Evaluation (OPE) in RL [58], as well as offline RL [59], is a related line of research, aiming to minimize the need for direct interaction with the environment to train or evaluate an RL agent. We use an instantiation of OPE and interpret it as a sampling mechanism [60] in our proposed methods. OPE approaches are also known to suffer from high variance in high-dimensional and complex settings. In general, RL and offline RL-based methods have yet to be adapted to complex environments such as CARLA, i.e., moving beyond tabular or simple settings often studied in OPE.
@@
Non-Reactive Simulation Benchmarks: Recent work by Dauner et al. [61] introduces NAVSIM, a data-driven non-reactive simulation framework that bridges offline and online evaluation by unrolling short-horizon simulations on real-world sensor data. It computes safety-critical metrics such as collisions and progress, using bird’s-eye-view abstractions. This approach has been shown to outperform traditional displacement errors (e.g., ADE) in closed-loop alignments. While NAVSIM enables scalable benchmarking (e.g., revealing lightweight models like TransFuser [41] can match complex architectures), its non-reactive assumption limits modeling of compounding errors or interactive scenarios. Our work complements NAVSIM by proposing uncertainty-weighted metrics that address rare or uncertain failures without requiring perception annotations, improving scalability in reactive or long-horizon settings.
@@ III-A Settings and Data
To validate model performance, we follow standard protocols in vision-based, end-to-end motion planners and open-source simulations. In our settings, we assume access to a collected dataset $\mathcal{D}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}_{i=1}^{N}$. Each observation $\mathbf{o}_{i}=(\mathbf{I}_{i},\mathbf{L}_{i},c_{i},v_{i})\in\mathcal{O}$, comprises an image $\mathbf{I}_{i}\in\mathbb{R}^{N_{W}\times N_{H}\times 3}$, a LiDAR point cloud with alpha channel $\mathbf{L}_{i}\in\mathbb{R}^{N_{X}\times N_{Y}\times N_{Z}\times 1}$, conditional command $c_{i}\in\mathbb{N}$, and speed measurements $v_{i}\in\mathbb{R}$, and ground-truth action vector $\mathbf{a}_{i}\in\mathcal{A}$ executed by the driver. We denote $\pi_{\boldsymbol{\theta}}:\mathcal{O}\to\mathcal{A}$ as the policy, a mapping function, parameterized using weights $\boldsymbol{\theta}\in\mathbb{R}^{d}$, that predicts actions $\hat{\mathbf{a}}_{i}$ based on input observations $\mathbf{o}_{i}$. We note that the action may either be a low-level control action, e.g., a 3D vector of steering, throttle, and brake amounts, or a set of future 2D waypoints in the map’s view representing an intended trajectory to follow by the vehicle [28, 46]. Vision-based planners today generally employ waypoints as the model output [10, 7, 41], where a low-level controller (e.g., PID) can be used to produce the final control. In contrast, prior work emphasizes direct steer value prediction [28]. Given the widespread use of waypoints-based metrics (defined below), our analysis explores the role of both action definitions and their correlation with online, closed-loop performance.
@@ III-B Metrics
A choice of setting weight scalars to $\alpha_{i}=1$ and setting an L1 loss, i.e., $\mathcal{L}=\|{\mathbf{a}_{i}-\mathbf{\hat{a}}_{i}}\|_{1}$, gives the mean absolute error (MAE) metric. Similarly, mean squared error (MSE) is given by using the L2 loss [7, 27]. In principle, a large number of different offline metrics may be derived for a given dataset, yet only a few have been shown to produce a correlation with online performance, particularly for policy’s overall success rate along a route known as the QCE and TRE (based on Codevilla et al. [28]), defined next for clarity. For clarity, we summarize all studied metrics in Table I.
@@
To ensure meaningful comparison, we employ the same threshold settings provided in Codevilla et al. [28].
@@
where $H$ is the Heaviside step function. We set $\lambda=0.1$ consistently with [28] and $\sigma=0.5$ as the midpoint of our steering range. We note that we set $\alpha_{i}=1$ in general for computing errors such as MSE, QCE, and TRE. However, we hypothesize that online errors occur in scenarios where the model may be uncertain about its outputs. We thus propose to leverage the weights $\alpha_{i}$ to proportionally weigh such instances, as defined next.
Uncertainty-Weighted Error (UWE): Codevilla et al. [28] were motivated by the intuition that errors at higher ego speeds have a greater impact on driving performance. They studied a speed-weighted metric where the weights in Eq. 1 are set to the speed, i.e., $\alpha_{i}=v_{i}$. However, this did not result in a better correlation with the route’s success rate. Instead, we propose to leverage a model-dependent, uncertainty-based weight. We employ monte carlo (MC) dropout [62] to efficiently estimate epistemic uncertainty. This involves enabling dropout during the prediction phase and performing $K$ forward passes through the network for each input sample $i$. We then calculate the variance $u_{i}$ of these predictions to estimate the model’s uncertainty [63]. Then, $u_{i}$ is raised to $\gamma$ to scale the variance’s importance, and a weighted sum is applied to various offline metrics, each with optimized weights $\beta_{t}$ to compute the UWE:
@@
where $n$ is the number of offline metrics and $\mathcal{L}_{t}$ is different offline metric. A weighted sum of the different uncertainty-based weighted offline metrics is able to capture various edge cases from each offline metric. While uncertainty has been applied to improve and regularize a range of robot perception and learning tasks [64, 37, 65, 66], we are the first to measure its utility in the context of offline and online policy evaluation. The MC Dropout-based uncertainty provides a simple and efficient weight that can be applied to any prior offline metric, yet we find its combinations with the basic errors, i.e., MAE and MSE, sufficient.
Online Metrics: Online metrics convey information about the driving performance in on-policy test time settings. We follow metrics defined in CARLA Leaderboard [45, 41], including the rate of various traffic incidences, such as Collisions and Stop Infractions. The main metric in CARLA is the Driving Score (DS). DS is the average of the product of the Route Completion (% from the planned route) score $R_{i}$ and an Infraction penalty $P_{i}$ across $N$ routes:
@@
We emphasize that prior work has only looked into Success Rate (SR) as the main online metric [28]. However, SR does not consider route completion (e.g., $95\%$ completion score still means zero success) nor infraction types along the route. Instead, DS provides a holistic evaluation of driving behavior, promoting safer and more efficient driving practices. In this paper, we will study the correlation of offline metrics with various online metrics, with a particular focus on the DS.

### v2

@@ Abstract
Real-world evaluation of perception-based planning models for robotic systems, such as autonomous vehicles, can be safely and inexpensively conducted offline, i.e. by computing model prediction error over a pre-collected validation dataset with ground-truth annotations. However, extrapolating from offline model performance to online settings remains a challenge. In these settings, seemingly minor errors can compound and result in test-time infractions or collisions. This relationship is understudied, particularly across diverse closed-loop metrics and complex urban maneuvers. In this work, we revisit this undervalued question in policy evaluation through an extensive set of experiments across diverse conditions and metrics. Based on analysis in simulation, we find an even worse correlation between offline and online settings than reported by prior studies, casting doubts on the validity of current evaluation practices and metrics for driving policies. Next, we bridge the gap between offline and online evaluation. We investigate an offline metric based on epistemic uncertainty, which aims to capture events that are likely to cause errors in closed-loop settings. The resulting metric achieves over 13% improvement in correlation compared to previous offline metrics. We further validate the generalization of our findings beyond the simulation environment in real-world settings, where even greater gains are observed.
@@ I Introduction
Can we estimate the online performance of a driving model without actively driving with it? Evaluating driving policy models that map camera input to a driving plan or controller action in real-world conditions is typically expensive and potentially hazardous. During model deployment, traffic infractions and collisions may occur, such as fatal accidents with pedestrians and vehicles [uber, cruiseend], Waymo’s driving in the wrong lane and veering unsafely [waymodrunk], and Tesla’s Autopilot crashing into a truck on a busy highway [ghorai2024causation].
To circumvent such safety-critical evaluations, researchers often use an offline approach. They leverage a computed expected loss over a pre-collected, real-world dataset. This method avoids the potentially dangerous and costly quantification of collisions and traffic infractions in live scenarios (Fig. 1). The efficiency in offline evaluation over a pre-collected driving dataset, where the driving model predictions are compared against safe ground-truth targets provided by a human driver (e.g., leveraging an L2 loss [hu2021safe, behl2020label, hu2023planning, zhang2023coaching, jaeger2023hidden, jiang2023vad, wang2019monocular, Codevilla2018ICRA, zhu2023learning]), has recently propelled significant advancements in scalable autonomous driving models [bansal2018chauffeurnet, hu2023planning, jiang2023vad, wu2023policy, dauner2023parting, hu2022st, zhang2024feedback, zhang2022rethinking]. However, it remains unclear whether current offline evaluation practices adequately reflect actual on-road driving model performance, particularly for decisions that require highly robust and split-second reactions with life-threatening implications.
@@
The reasons for the discrepancy between offline and online model performance are three-fold. First, real-world driving datasets are nearly entirely uneventful, with drivers rarely experiencing safety-critical events such as abrupt pedestrian crossings, near-collisions, or even turns and merging. The expected loss over the inherently imbalanced distribution of data can obscure performance under safety-critical events [tu2023towards, li2023ego]. Second, while automatic data sampling techniques can alleviate data imbalance issues [thorn2018framework, menzel2018scenarios, o2018scalable, weber2019framework, lee2020adaptive], the standard L2 loss-based metric itself provides a crude measure for the ultimate task of safe and enjoyable driving [li2020end, Codevilla2018ECCV]. For instance, the L2 metric can remain identical among several maneuvers, e.g., merging to the left or right lane on a freeway, yet one decision may be perfectly safe while the other results in a collision, i.e., in the case of an adjacent vehicle in a nearby lane. This observation holds for nearly any scenario in complex urban driving settings, as shown in Fig. 1. Additional open-loop metrics have been devised, such as those based on future collision estimates [hu2021safe, hu2023planning, jaeger2023hidden, jiang2023vad, li2023ego]. However, these metrics induce similar evaluation issues. Their reliance on various naive velocity models prevents them from capturing realistic interactive navigation and negotiation in intricate and dense settings. Moreover, collision metrics lack scalability due to the requirement of perception annotations [schreier2023offline]. Third, online model testing is done in a closed-loop manner, where minor decision errors can propagate and result in unpredictable model behavior over time in testing [ross2011reduction, osa2018algorithmic, prakash2020exploring, laskey2017dart].
@@
Despite known limitations in offline open-loop evaluation [zhang2022rethinking, li2023ego], prior work has rarely studied extrapolation from offline to online settings beyond simplistic settings. For instance, Codevilla et al. [Codevilla2018ECCV] demonstrated the limited correlation between the two settings but only in the simplest towns and maneuvers in the CARLA simulation [Dosovitskiy2017CORL], i.e., Town01 and Town02, with one online metric (success rate), and partial action values (only steer, excluding throttle and brake ground truth). Therefore, generalization to intricate maneuvers and complex urban scenarios, diverse online metrics, or real-world settings remains understudied despite its broad implications. In this work, we revisit the question of the correlation between offline and online evaluations.
@@
Contributions: We aim to advance safe and reliable evaluation procedures for autonomous driving policies. Our key insights are threefold. First, we leverage a more expansive evaluation in simulation, covering multiple towns and closed-loops metrics. Our results reveal even worse correlation patterns than have been previously shown. Subsequently, we hypothesize that the poor correlation is due to degraded model performance during instances of uncertainty, i.e., challenging or rare events in the data. We thus analyze an effective metric based on an uncertainty-weighted error, and demonstrate its broad effectiveness across settings, models, and error types. While uncertainty-based techniques are standard in robotics and machine learning [filos2020can, stachowicz2024racer, ovadia2019can], to the best of our knowledge, we are the first to quantify their role in estimating online evaluation of driving policies. We observe a significant improvement, e.g., 13% higher correlation in CARLA. Third, given challenges in realistic simulation of real-world physics and appearance, we go beyond simplistic simulations to characterize correlation trends in the real world using a small-scale platform, further validating our proposed metric.
@@ II Related Work
End-to-End Autonomous Driving: With the emergence of deep learning models, autonomous driving research has undergone a significant transformation, shifting from rule-based systems to learning-based approaches. This shift is notably reflected in the proliferation of end-to-end autonomous driving policies utilizing deep learning techniques [bojarski2016end, lai2024uncertainty, xu2017end, chitta2022transfuser, zhang2023coaching, pomerleau1989alvinn, bansal2018chauffeurnet, Cui_2021_ICCV, behl2020label, ohn2020learning, hu2022st]. This trend is further exemplified by the CARLA Leaderboard [carlaleaderboard], where deep learning-based methods have gained prominence, surpassing modular approaches in autonomous driving performance [muller2018driving, tang2018teaching, chen2021what, gonzalez2015review, kendall2019learning, xu2021autonomous]. In addition to camera-based RGB images, additional sensors, such as LiDAR, have also been shown to play a key role in improving the safety of driving policies through multi-modal sensor fusion [chitta2022transfuser, shao2023safety, prakash2021multi].
@@
Issues with Offline Evaluation: The progress of learning-based models is primarily driven by offline evaluation methods. However, the question of their accuracy in assessing driving performance within online settings still remains. Codevilla et al. [Codevilla2018ECCV] is one of the first to attempt to study this problem and proposes several novel offline evaluation metrics, including speed-weighted and quantized versions of standard loss metrics. Among the metrics, the NuScenes Detection Score (NDS) [caesar2020nuscenes] is a widely accepted offline metric that has been shown in a recent work by Schreier et al. [schreier2023offline] to provide improved correlation with online metrics like CARLA Driving Score [Dosovitskiy2017CORL]. As shown by Li et al. in [li2023ego], a simple baseline model without vision can achieve competitive status on a leaderboard evaluated by offline metrics like NDS, thus proving the limitations of existing offline metrics in predicting driving performance. These recent discussions demand a revisit on whether offline metrics are sufficient in predicting driving performance.
@@
Safety-Critical Scenario Analysis: Safety critical scenarios refer to specific scenarios, such as collisions, where safety is compromised, and the ego vehicle must take preventative or evasive actions to avoid catastrophic outcomes [ghorai2024causation, wang2021advsim, weber2019framework, hanselmann2022king]. Current methodologies address this challenge by curating datasets with generated safety-critical scenarios [lee2020adaptive, lai2023xvo, menzel2018scenarios, thorn2018framework]. This approach allows for testing on various generated scenarios to validate if the policy is safe. While we share a similar goal, this is an orthogonal direction to our approach as we focus on designing a useful metric across datasets. However, while generated scenario realism and coverage remain an open question, our proposed metric can be automatically and easily applied at scale to any policy and dataset (i.e., only based on the model and without requiring privileged annotated information).
@@
Off-Policy Evaluation: Offline Policy Evaluation (OPE) in RL [review-off-RL-tutorial], as well as offline RL [off-RL-tutorial], is a related line of research, aiming to minimize the need for direct interaction with the environment to train or evaluate an RL agent. We use an instantiation of OPE and interpret it as a sampling mechanism [instantiation-uncertainty-based-metric] in our proposed methods. OPE approaches are also known to suffer from high variance in high-dimensional and complex settings. In general, RL and offline RL-based methods have yet to be adapted to complex environments such as CARLA, i.e., moving beyond tabular or simple settings often studied in OPE.
@@
Non-Reactive Simulation Benchmarks: Recent work by Dauner et al. [dauner2024navsim] introduces NAVSIM, a data-driven non-reactive simulation framework that bridges offline and online evaluation by unrolling short-horizon simulations on real-world sensor data. It computes safety-critical metrics such as collisions and progress, using bird’s-eye-view abstractions. This approach has been shown to outperform traditional displacement errors (e.g., ADE) in closed-loop alignments. While NAVSIM enables scalable benchmarking (e.g., revealing lightweight models like TransFuser [chitta2022transfuser] can match complex architectures), its non-reactive assumption limits modeling of compounding errors or interactive scenarios. Our work complements NAVSIM by proposing uncertainty-weighted metrics that address rare or uncertain failures without requiring perception annotations, improving scalability in reactive or long-horizon settings.
@@ III-A Settings and Data
To validate model performance, we follow standard protocols in vision-based, end-to-end motion planners and open-source simulations. In our settings, we assume access to a collected dataset $\mathcal{D}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}_{i=1}^{N}$. Each observation $\mathbf{o}_{i}=(\mathbf{I}_{i},\mathbf{L}_{i},c_{i},v_{i})\in\mathcal{O}$, comprises an image $\mathbf{I}_{i}\in\mathbb{R}^{N_{W}\times N_{H}\times 3}$, a LiDAR point cloud with alpha channel $\mathbf{L}_{i}\in\mathbb{R}^{N_{X}\times N_{Y}\times N_{Z}\times 1}$, conditional command $c_{i}\in\mathbb{N}$, and speed measurements $v_{i}\in\mathbb{R}$, and ground-truth action vector $\mathbf{a}_{i}\in\mathcal{A}$ executed by the driver. We denote $\pi_{\boldsymbol{\theta}}:\mathcal{O}\to\mathcal{A}$ as the policy, a mapping function, parameterized using weights $\boldsymbol{\theta}\in\mathbb{R}^{d}$, that predicts actions $\hat{\mathbf{a}}_{i}$ based on input observations $\mathbf{o}_{i}$. We note that the action may either be a low-level control action, e.g., a 3D vector of steering, throttle, and brake amounts, or a set of future 2D waypoints in the map’s view representing an intended trajectory to follow by the vehicle [Codevilla2018ECCV, muller2018driving]. Vision-based planners today generally employ waypoints as the model output [jiang2023vad, hu2023planning, chitta2022transfuser], where a low-level controller (e.g., PID) can be used to produce the final control. In contrast, prior work emphasizes direct steer value prediction [Codevilla2018ECCV]. Given the widespread use of waypoints-based metrics (defined below), our analysis explores the role of both action definitions and their correlation with online, closed-loop performance.
@@ III-B Metrics
A choice of setting weight scalars to $\alpha_{i}=1$ and setting an L1 loss, i.e., $\mathcal{L}=\|{\mathbf{a}_{i}-\mathbf{\hat{a}}_{i}}\|_{1}$, gives the mean absolute error (MAE) metric. Similarly, mean squared error (MSE) is given by using the L2 loss [hu2023planning, li2020end]. In principle, a large number of different offline metrics may be derived for a given dataset, yet only a few have been shown to produce a correlation with online performance, particularly for policy’s overall success rate along a route known as the QCE and TRE (based on Codevilla et al. [Codevilla2018ECCV]), defined next for clarity. For clarity, we summarize all studied metrics in Table I.
@@
To ensure meaningful comparison, we employ the same threshold settings provided in Codevilla et al. [Codevilla2018ECCV].
@@
where $H$ is the Heaviside step function. We set $\lambda=0.1$ consistently with [Codevilla2018ECCV] and $\sigma=0.5$ as the midpoint of our steering range. We note that we set $\alpha_{i}=1$ in general for computing errors such as MSE, QCE, and TRE. However, we hypothesize that online errors occur in scenarios where the model may be uncertain about its outputs. We thus propose to leverage the weights $\alpha_{i}$ to proportionally weigh such instances, as defined next.
Uncertainty-Weighted Error (UWE): Codevilla et al. [Codevilla2018ECCV] were motivated by the intuition that errors at higher ego speeds have a greater impact on driving performance. They studied a speed-weighted metric where the weights in Eq. 1 are set to the speed, i.e., $\alpha_{i}=v_{i}$. However, this did not result in a better correlation with the route’s success rate. Instead, we propose to leverage a model-dependent, uncertainty-based weight. We employ monte carlo (MC) dropout [gal2016dropout] to efficiently estimate epistemic uncertainty. This involves enabling dropout during the prediction phase and performing $K$ forward passes through the network for each input sample $i$. We then calculate the variance $u_{i}$ of these predictions to estimate the model’s uncertainty [lai2025zerovo]. Then, $u_{i}$ is raised to $\gamma$ to scale the variance’s importance, and a weighted sum is applied to various offline metrics, each with optimized weights $\beta_{t}$ to compute the UWE:
@@
where $n$ is the number of offline metrics and $\mathcal{L}_{t}$ is different offline metric. A weighted sum of the different uncertainty-based weighted offline metrics is able to capture various edge cases from each offline metric. While uncertainty has been applied to improve and regularize a range of robot perception and learning tasks [kendall2018multi, ovadia2019can, henaff2019model, wang2021uncertainty], we are the first to measure its utility in the context of offline and online policy evaluation. The MC Dropout-based uncertainty provides a simple and efficient weight that can be applied to any prior offline metric, yet we find its combinations with the basic errors, i.e., MAE and MSE, sufficient.
Online Metrics: Online metrics convey information about the driving performance in on-policy test time settings. We follow metrics defined in CARLA Leaderboard [carlaleaderboard, chitta2022transfuser], including the rate of various traffic incidences, such as Collisions and Stop Infractions. The main metric in CARLA is the Driving Score (DS). DS is the average of the product of the Route Completion (% from the planned route) score $R_{i}$ and an Infraction penalty $P_{i}$ across $N$ routes:
@@
We emphasize that prior work has only looked into Success Rate (SR) as the main online metric [Codevilla2018ECCV]. However, SR does not consider route completion (e.g., $95\%$ completion score still means zero success) nor infraction types along the route. Instead, DS provides a holistic evaluation of driving behavior, promoting safer and more efficient driving practices. In this paper, we will study the correlation of offline metrics with various online metrics, with a particular focus on the DS.